OpenStudy (anonymous):

Find the total area for the regular pyramid. Sides: 12 and 10

jhonyy9 (jhonyy9):

base area =12*12=144 height of a triangular side =sqrt(100-36)=sqrt64 =8 area of a side=12*8/2 =12*4 =48 lateral area =4*48=192 total area =144 + 192 =336